# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 3. $(C,\alpha)$-good property
Case 2.2. $a(s)$ is unbounded. Here $a(s)\asymp s^{\theta}$, for $\theta>0$. We recall that for each $\mathrm{n}\in\mathbb{N}$ there is $T_{\mathrm{n}}$ such that all basis functions $F^{-}_{i},F_{j}^{+}$ will be continuously differentiable $\mathrm{n}$ times in the ray $[T_{\mathrm{n}},\infty)$ for all $i,j$, see [Geom_cat_and_o_min_str]. By [Miller94], it holds that
$\left|\frac{d^{p}}{dt^{p}}F^{-}_{i}(x)\right|\ll x^{-\kappa_{i}-p},\text{ and }\left|\frac{d^{q}}{dt^{q}}F^{+}_{j}(x)\right|\ll x^{r_{j}-q},$
where $p,q$ are non-negative integers.
Let $J(s)=[a(s),a(s)+l(s)]$. Then we conclude that for an integer $\mathrm{n}$ such that $r_{i}-(\mathrm{n}+1)<0$ for all $0\leq i\leq m$, we have
$\sup_{t\in J(s)}\left\lvert\frac{d^{\mathrm{n}+1}}{dt^{\mathrm{n}+1}}f_{\hat{\underline{c}}(s)}(t)\right\rvert=O(a(s)^{r_{m}-(\mathrm{n}+1)}).$
By Taylor’s theorem, for the Taylor polynomial $P_{\mathrm{n},a(s)}(t)$ of $f_{\hat{\underline{c}}(s)}(t)$ of degree $\mathrm{n}$ centered around $a(s)$, we have
$\sup_{t\in J(s)}|f_{\hat{\underline{c}}(s)}(t)-P_{\mathrm{n},a(s)}(t)|$ $=O(a(s)^{r_{m}-(\mathrm{n}+1)}l(s)^{\mathrm{n}+1})$
$=O\left(a(s)^{r_{m}}\left(\frac{l(s)}{a(s)}\right)^{\mathrm{n}+1}\right)$
(3.27) $=O\left(s^{r_{m}\theta}s^{-(\mathrm{n}+1)\kappa}\right),$
where we use the fact that $a(s)/l(s)\sim C_{0}s^{\kappa}$ for some $C_{0}\neq 0$ and $\kappa>0$. Due to Lemma 3.5, the function $\Psi(s):=\|f_{\hat{\underline{c}}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}$ is positive. Since $\Psi$ is also definable in a polynomially bounded o-minimal structure, we pick $\eta>0$ such that
(3.28) $\|f_{\hat{\underline{c}}(s)}\|_{J(s)}\gg s^{-\eta},~{}\text{as }s\to\infty.$
We take $\mathrm{n}$ large enough such that
$\nu:=r_{m}\theta-(\mathrm{n}+1)\kappa<-\eta.$
Then, for $J_{\delta}(s)=[\alpha(s),\alpha(s)+\delta l(s)]$,
(by (3.27)) $\|f_{\hat{\underline{c}}(s)}\|_{J(s)}\leq$ $\|P_{\mathrm{n},a(s)}\|_{J(s)}+O(s^{\nu})$
(by Proposition 3.4) $\leq$ $M_{\mathrm{n}}(\delta)\|P_{\mathrm{n},a(s)}\|_{J_{\delta}(s)}+O(s^{\nu})$
(by (3.27)) $\leq$ $M_{\mathrm{n}}(\delta)\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}+O(s^{\nu})$
$=$ $M_{\mathrm{n}}(\delta)(\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}\left[1+O\left(\frac{s^{\nu}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}\right)\right]$
(by (3.28)) $\ll$ $M_{\mathrm{n}}(\delta)\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}\left[1+O\left(\frac{s^{\nu}}{s^{-\eta}}\right)\right]$
Namely, $\frac{\|f_{\hat{\underline{c}}(s)}\|_{J(s)}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}$ is bounded for all large $s$, contradicting assumption $s<\frac{\|f_{\hat{\underline{c}}(s)}\|_{J(s)}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}$. ∎
@@ 3.3. Relative time near varieties
Using the $(C,\alpha)$-good property, we prove the following proposition. It is an analog of [Dani1993LimitDO, Proposition 4.2], and it is key for the linearization technique.
@@
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous, unbounded curve definable in a polynomially bounded o-minimal structure such that $\lim_{t\to\infty}\psi(t)v\neq 0,\forall v\in\mathbb{R}^{m}\smallsetminus\{0\}$. Consider a non-zero polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$, and let
$\emph{\text{V}}:=\{x\in\mathbb{R}^{m}:Q(x)=0\}.$
Then, there exists $T_{0}>0$ such that the following holds: for an $\epsilon>0$ and a compact subset $\emph{\text{K}}_{1}\subset\emph{\text{V}}$, there exists a compact subset $\emph{\text{K}}_{2}$ with
$\emph{\text{K}}_{1}\subset\emph{\text{K}}_{2}\subset\emph{\text{V}}$
such that for every compact neighborhood $\Phi$ of $K_{2}$ in $\mathbb{R}^{m}$, there exists a compact neighborhood $\Psi$ of $K_{1}$ in $\mathbb{R}^{m}$, with $\Psi\subset\mathring{\Phi}$, where $\mathring{\Phi}$ denotes the interior of $\Phi$, such that:
(3.29) |$|\{t\in[a,b]:\psi(t)v\in\Psi\}|\leq\epsilon(b-a).$
@@
Compared to [Dani1993LimitDO, Proposition 4.2], the statement of the above proposition is different in the assumption that the curve $\psi(t)v$ remains in $\overline{\Phi}$ throughout the whole interval $[a,b]$, and exists $\Phi$ at $t=b$. This is a stronger requirement than the one in [Dani1993LimitDO, Proposition 4.2] which only asks for $\psi(t_{0})v\notin\Phi$ for some $t_{0}\in[a,b]$.
@@
Notice that for any affine variety $\text{V}\subset\mathbb{R}^{m}$ there exists a polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$ such that V is the zero set of $Q$. Indeed, by the Hilbert basis theorem, V is the zero set of finitely many polynomials $f_{1},...,f_{r}$, and we put $Q:=f_{1}^{2}+\cdots+f_{r}^{2}$.
@@
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous curve definable in a polynomially bounded o-minimal structure, and let $Q:\mathbb{R}^{m}\to\mathbb{R}$ be a polynomial. Then there exist constants $\nu,C,\alpha,T_{0}>0$ such that for any $v\in\mathbb{R}^{m}$ it holds that either:
@@
Consider the vector space $\mathcal{V}$, of functions from $[0,\infty)$ to $\mathbb{R}$, defined by:
$\mathcal{V}=\text{Span}_{\mathbb{R}}\{t\mapsto Q(\psi(t)v):v\in\mathbb{R}^{m}\}.$
It is straightforward to verify that $\mathcal{V}$ is finite dimensional, and that all functions in $\mathcal{V}$ are definable. In view of Remark 3.8, there is a $T_{0}\geq 0$ such that if $f\in\mathcal{V}$ is eventually zero, then $f(t)=0,\forall t\in[T_{0},\infty)$. We consider $\mathscr{F}$ to be the space of functions of $\mathcal{V}$ restricted to $[T_{0},\infty).$ By Lemma 3.7 that there is a basis $\{f_{0},f_{1},...,f_{n}\}$ for $\mathscr{F}$ such that $\deg(f_{0})<\deg(f_{1})<...<\deg(f_{n})$. Let $\nu>0$ such that $\deg(t^{\nu}f_{0})\geq 0$. Then, by Theorem 1.14, we pick $T_{1}>0$ such that any non-zero function
$f\in t^{\nu}\otimes\mathscr{F}=\text{Span}_{\mathbb{R}}\{t\mapsto t^{\nu}Q(\psi(t)v):v\in\mathbb{R}^{m}\}$
is $(C,\alpha)$-good in $[T_{1},\infty)$. ∎
@@
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous curve, definable in a polynomially bounded o-minimal structure. Fix a polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$. Then there exist $T_{0},C,\alpha>0$ such that for any $v\in\mathbb{R}^{m}$ and any interval $[x,y]\subseteq[T_{0},\infty)$, if $\sup_{t\in[x,y]}|Q(\psi(t)v)|=|Q(\psi(y)v)|$, then
@@
Let $\nu,C,\alpha,T_{0}>0$ be as in Proposition 3.13. Let $[x,y]\subseteq[T_{0},\infty)$ and consider a vector $v\in\mathbb{R}^{m}$ with $|Q(\psi(y)v)|=\delta>0$ and $|Q(\psi(t)v)|\leq\delta$ for all $t\in[x,y]$. Then it follows that
@@
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous unbounded curve definable in a polynomially bounded o-minimal structure such that $\lim_{t\to\infty}\psi(t)v\neq 0,\forall v\neq 0$. Pick $T_{0},C,\alpha>0$ such that for all $v\in\mathbb{R}^{m}$, the $(C,\alpha)$-good property holds for the norm-map $t\mapsto\lVert\psi(t)v\rVert$ as in Proposition 1.15, and the right-max-$(C,\alpha)$-property holds for the map $t\mapsto\lvert Q(\psi(t)v)\rvert$ as in Corollary 3.14. Let $\epsilon>0$ be given.
Given a compact set $\text{K}_{1}\subset\text{V}$, let $R_{1}>0$ be such that $\text{K}_{1}\subset B_{R_{1}}(0)$. Let $\text{K}_{2}:=\overline{\text{B}_{R_{2}}(0)}\cap\text{V}$, where $R_{2}>R_{1}$ is such that $C(R_{1}/R_{2})^{\alpha}\leq\epsilon$, see (3.34).
For any $\epsilon^{\prime}>0$, define
$\text{V}^{\epsilon^{\prime}}:=\{x\in\mathbb{R}^{m}:|Q(x)|<\epsilon^{\prime}\}.$
@@
(3.32) $B_{R_{2}}(0)\cap\text{V}^{\epsilon_{1}}\subset\mathring{\Phi}.$
We pick $\epsilon_{2}\in(0,\epsilon_{1})$ such that $K_{1}\subset B_{R_{1}}(0)\cap\text{V}^{\epsilon_{2}}$ and $C(\epsilon_{2}/\epsilon_{1})^{\alpha}\leq\epsilon$, see (3.35). Let
(3.33) $\Psi:=\overline{B_{R_{1}}(0)\cap\text{V}^{\epsilon_{2}}}.$
Let $v\in\mathbb{R}^{m}$ and an interval $[a,b]\subset[T_{0},\infty)$ such that:
@@
  • $\psi(b)v\in\Phi\smallsetminus\mathring{\Phi}$.
Then by (3.32), $\|\psi(b)v\|\geq R_{2}$ or $|Q(\psi(b)v)|\geq\epsilon_{1}.$ Now, if $\|\psi(b)v\|\geq R_{2}$, then by (3.33) and Proposition 1.15,
@@
Suppose now that $|Q(\psi(b)v)|\geq\epsilon_{1}$. Since $t\mapsto|Q(\psi(t)v)|$ is continuous and definable, and since a definable map takes a given value only finitely many times, there is a decomposition
@@
for some $k\in\mathbb{N}$, where for all $1\leq i\leq k$ it holds
@@
Now, since $\Psi=B_{R_{1}}(0)\cap\text{V}^{\epsilon_{2}}$, we have
$\{t\in[a,b]|\psi(t)v\in\Psi\}\subseteq\{t\in[a,b]:|Q(\psi(t)v)|\leq\epsilon_{2}\},$
@@
for each $1\leq i\leq k$. Thus we may conclude that

### v2

@@ 3. $(C,\alpha)$-good property
Case 2.2. $a(s)$ is unbounded. Here $a(s)\asymp s^{\theta}$, for $\theta>0$. We recall that for each $\mathrm{n}\in\mathbb{N}$, there is a $T_{\mathrm{n}}$ such that all basis functions $F^{-}_{i},F_{j}^{+}$ will be continuously differentiable $\mathrm{n}$ times in the ray $[T_{\mathrm{n}},\infty)$ for all $i,j$ (see Theorem B.5). By Lemma 3.6, it holds that
$\left|\frac{d^{p}}{dt^{p}}F^{-}_{i}(x)\right|\ll_{p}x^{-\kappa_{i}-p},\text{ and }\left|\frac{d^{q}}{dt^{q}}F^{+}_{j}(x)\right|\ll_{q}x^{r_{j}-q},$
where $\kappa_{i},r_{j}$ are as in (3.5), $p,q$ are non-negative integers, and the implied constants depend on $p,q$. Then, for all sufficiently large integers $\mathrm{n}$ such that $r_{i}-(\mathrm{n}+1)<0$ for all $0\leq i\leq m$, it holds that
$\sup_{t\in[a(s),a(s)+l(s)]}\left\lvert\frac{d^{\mathrm{n}+1}}{dt^{\mathrm{n}+1}}f_{\hat{\underline{c}}(s)}(t)\right\rvert=O_{\mathrm{n}}(a(s)^{r_{m}-(\mathrm{n}+1)}),$
where the implied constant depends on $\mathrm{n}$. By Taylor’s theorem, for the Taylor polynomial $P_{\mathrm{n},a(s)}(t)$ of $f_{\hat{\underline{c}}(s)}(t)$ of degree $\mathrm{n}$ centered around $a(s)$, we have
$\sup_{t\in[a(s),a(s)+l(s)]}|f_{\hat{\underline{c}}(s)}(t)-P_{\mathrm{n},a(s)}(t)|$ $=O_{\mathrm{n}}(a(s)^{r_{m}-(\mathrm{n}+1)}l(s)^{\mathrm{n}+1})$
$=O_{\mathrm{n}}\left(a(s)^{r_{m}}\left(\frac{l(s)}{a(s)}\right)^{\mathrm{n}+1}\right)$
(3.25) $=O_{\mathrm{n}}\left(s^{r_{m}\theta}s^{-(\mathrm{n}+1)\kappa}\right),$
where the last line follows by the assumption that $a(s)/l(s)\sim s^{\kappa}$ with $\kappa>0$. Due to Lemma 3.7, the function $\Psi(s):=\|f_{\hat{\underline{c}}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}$ is positive. Since $\Psi$ is definable in a polynomially bounded o-minimal structure, there exists an $\eta>0$ such that
(3.26) $\|f_{\hat{\underline{c}}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}\gg s^{-\eta},~{}\text{as }s\to\infty.$
We fix $\mathrm{n}$ large enough such that $\nu:=r_{m}\theta-(\mathrm{n}+1)\kappa<-\eta.$ We denote $J(s)=[a(s),a(s)+l(s)]$ and we let $J_{\delta}(s)=[\alpha(s),\alpha(s)+\delta l(s)]$. Then:
[…]
$=M$ (${}_{\mathrm{n}}(\delta)\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}\left[1+O\left(\frac{s^{\nu}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}\right)\right]$
[…]
Namely, $\frac{\|f_{\hat{\underline{c}}(s)}\|_{J(s)}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}$ is bounded for all large $s$, contradicting the assumption $s<\frac{\|f_{\hat{\underline{c}}(s)}\|_{J(s)}}{\|f_{\hat{\underline{c}}(s)}\|_{J_{\delta}(s)}}$. ∎
@@ 3.3. Relative time near varieties
The following proposition is an analog of [Dani1993LimitDO, Proposition 4.2], and it is key for the linearization technique.
@@
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous, unbounded curve definable in $\mathscr{S}_{\mathrm{pbd}}$ such that $\lim_{t\to\infty}\psi(t)v\neq 0,\forall v\in\mathbb{R}^{m}\setminus\{0\}$. Fix a nonzero polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$, and denote
$\mathcal{A}:=\{x\in\mathbb{R}^{m}:Q(x)=0\}.$
Then, there exists a $T_{0}>0$ such that the following holds: for an $\epsilon>0$ and a compact subset $\emph{\text{K}}_{1}\subseteq\mathcal{A}$, there exists a compact subset $\emph{\text{K}}_{2}$ with
$\emph{\text{K}}_{1}\subseteq\emph{\text{K}}_{2}\subseteq\mathcal{A}$
such that for every compact neighborhood $\Phi$ of $K_{2}$ in $\mathbb{R}^{m}$, there exists a compact neighborhood $\Psi$ of $K_{1}$ in $\mathbb{R}^{m}$, with $\Psi\subseteq\mathring{\Phi}$, where $\mathring{\Phi}$ denotes the interior of $\Phi$, such that:
(3.27) |$|\{t\in[a,b]:\psi(t)v\in\Psi\}|\leq\epsilon(b-a),$
@@
Compared to [Dani1993LimitDO, Proposition 4.2], the statement our proposition is different in the assumption that the curve $\psi(t)v$ remains in $\Phi$ throughout the whole interval $[a,b]$, and exits $\mathring{\Phi}$ at $t=b$. This is a stronger requirement than the one in [Dani1993LimitDO, Proposition 4.2] which only asks for $\psi(t_{0})v\notin\mathring{\Phi}$ for some $t_{0}\in[a,b]$.
@@
Notice that for any affine variety $\mathcal{A}\subseteq\mathbb{R}^{m}$ there exists a polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$ such that $\mathcal{A}$ is the zero set of $Q$. Indeed, by the Hilbert basis theorem, $\mathcal{A}$ is the zero set of finitely many polynomials $f_{1},...,f_{r}$, and we let $Q:=f_{1}^{2}+\cdots+f_{r}^{2}$.
@@
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous curve definable in $\mathscr{S}_{\mathrm{pbd}}$, and let $Q:\mathbb{R}^{m}\to\mathbb{R}$ be a polynomial. Then there exist constants $\nu,C,\alpha,T_{0}>0$ such that for any $v\in\mathbb{R}^{m}$ exactly one of the following properties hold:
@@
Consider the vector space
$\mathcal{V}:=\text{Span}_{\mathbb{R}}\{Q(\psi(t)v),\ t\geq 0:v\in\mathbb{R}^{m}\}.$
It is straightforward to verify that $\mathcal{V}$ is finite-dimensional, and that all functions in $\mathcal{V}$ are definable. In view of Remark 3.10, there is a $T_{0}\geq 0$ such that if $f\in\mathcal{V}$ is eventually zero, then $f(t)=0,\forall t\in[T_{0},\infty)$. We consider $\mathscr{F}$ to be the space of functions of $\mathcal{V}$ restricted to $[T_{0},\infty).$ By Lemma 3.9, there is a basis $\{f_{0},f_{1},...,f_{n}\}$ for $\mathscr{F}$ such that $\deg(f_{0})<\deg(f_{1})<...<\deg(f_{n})$. Let $\nu>0$ such that $\deg(t^{\nu}f_{0})\geq 0$. Then, by Theorem 1.12, we pick $T_{1}>0$ such that any nonzero function
$f\in t^{\nu}\otimes\mathscr{F}=\text{Span}_{\mathbb{R}}\{t^{\nu}Q(\psi(t)v),\ t\geq T_{1}:v\in\mathbb{R}^{m}\}$
is $(C,\alpha)$-good. ∎
@@
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a continuous curve, definable in $\mathscr{S}_{\mathrm{pbd}}$. Fix a polynomial $Q:\mathbb{R}^{m}\to\mathbb{R}$. Then there exist $T_{0},C,\alpha>0$ such that for any $v\in\mathbb{R}^{m}$ and any interval $[x,y]\subseteq[T_{0},\infty)$, if
$\sup_{t\in[x,y]}|Q(\psi(t)v)|=|Q(\psi(y)v)|,$
then
@@
Fix $\nu,C,\alpha,T_{0}>0$ as in Proposition 3.15. Let $[x,y]\subseteq[T_{0},\infty)$ and let $v\in\mathbb{R}^{m}$ be a vector such that
$|Q(\psi(y)v)|=\delta>0,$
and such that
$|Q(\psi(t)v)|\leq\delta,\ \forall t\in[x,y].$
Then,
@@
Let $\psi:[0,\infty)\to\text{GL}(m,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous unbounded curve such that $\lim_{t\to\infty}\psi(t)v\neq 0,\forall v\neq 0$. Pick $T_{0},C,\alpha>0$ such that the $(C,\alpha)$-good property holds for the family
$\mathscr{F}_{\text{norm}}:=\{\|\psi(t)v\|,\ t\geq T_{0}:v\in\mathbb{R}^{m}\setminus\{0\}\},$
see Proposition 1.13, and such that the right-max-$(C,\alpha)$-property holds for the functions in
$\mathscr{F}_{\text{variety}}:=\{\lvert Q(\psi(t)v)\rvert,\ t\geq T_{0}:v\in\mathbb{R}^{m}\},$
which satisfy the requirements of Corollary 3.16.
Let $\epsilon>0$. Given a compact set $\text{K}_{1}\subseteq\mathcal{A}$, let $R_{1}>0$ be such that $\text{K}_{1}\subseteq B_{R_{1}}(0)$. We choose $R_{2}>R_{1}$ is such that $C(R_{1}/R_{2})^{\alpha}\leq\epsilon$, and we denote $\text{K}_{2}:=\overline{\text{B}_{R_{2}}(0)}\cap\mathcal{A}$; see (3.32) for the use of this choice. We define for $\epsilon^{\prime}>0$:
$\mathcal{A}^{\epsilon^{\prime}}:=\{x\in\mathbb{R}^{m}:|Q(x)|<\epsilon^{\prime}\}.$
@@
(3.30) $B_{R_{2}}(0)\cap\mathcal{A}^{\epsilon_{1}}\subseteq\mathring{\Phi}.$
We denote $\epsilon_{2}:=\frac{R_{1}}{R_{2}}\epsilon_{1}$, and we note that $K_{1}\subseteq B_{R_{1}}(0)\cap\mathcal{A}^{\epsilon_{2}}$ and we observe that $C(\epsilon_{2}/\epsilon_{1})^{\alpha}\leq\epsilon$; see (3.33) for the use of this choice. We consider
(3.31) $\Psi:=\overline{B_{R_{1}}(0)\cap\mathcal{A}^{\epsilon_{2}}}.$
Let $v\in\mathbb{R}^{m}$ and an interval $[a,b]\subseteq[T_{0},\infty)$ such that:
@@
  • $\psi(b)v\in\Phi\setminus\mathring{\Phi}$.
Then by (3.30), $\|\psi(b)v\|\geq R_{2}$ or $|Q(\psi(b)v)|\geq\epsilon_{1}.$
Suppose that $\|\psi(b)v\|\geq R_{2}$. By (3.31) and by Proposition 1.13,
@@
Suppose now that $|Q(\psi(b)v)|\geq\epsilon_{1}$. Since $t\mapsto|Q(\psi(t)v)|$ is continuous and definable, and since a definable map takes a particular value only finitely many times, there is a decomposition
@@
for some $k\in\mathbb{N}$, such that for all $1\leq i\leq k$:
@@
Since $\Psi\subseteq\mathcal{A}^{\epsilon_{2}}$, we have
$\{t\in[a,b]:\psi(t)v\in\Psi\}\subseteq\{t\in[a,b]:|Q(\psi(t)v)|\leq\epsilon_{2}\},$
@@
for each $1\leq i\leq k$. This proves that